A body of mass 2 kg moving with a velocity of 5 ms-1 collides with a stationary body of mass 1/2 kg. If the two bodies move together after the impact, calculate their common velocity.
1
Expert's answer
2020-06-29T14:06:43-0400

We can find the common velocity of two bodies from the law of conservation of momentum:


"m_1v_1 + m_2v_2 = (m_1+m_2)v_{common},""v_{common} = \dfrac{m_1v_1 + m_2v_2}{m_1 + m_2},""v_{common} = \dfrac{2 \ kg \cdot 5 \ \dfrac{m}{s} + 0.5 \ kg \cdot 0 \ \dfrac{m}{s}}{2 \ kg + 0.5 \ kg} = 4 \ \dfrac{m}{s}."

Answer:

"v_{common} = 4 \ \dfrac{m}{s}."